# Phase diagram of the quantum spin-1/2 Heisenberg-$\Gamma$ model on a frustrated zigzag chain

### v1

@@ II.1 Model Hamiltonian
Figure 5: (a) Magnetization curves for field $h$ applied along the $z$-direction. We choose $(\phi,\theta)/\pi=(0.05,0.05)$ for TLL, $(0.2,0.05)$ for ND, $(0.75,0.1)$ for VC, and $(0.85,0.3)$ for N phases. Among them, only the ND phase has a spin gap $\Delta\approx 0.3$. (b) Nematic parameter $\overline{\langle Q^{xy}_{i}Q^{xy}_{j}\rangle}$ averaged over $|i-j|=58-61$ as a function of $\Gamma/J$ for the choices of $\phi/\pi$ corresponding to those of panel (a). All of them behave in power of $\Gamma/J$.
II.2 Ground-state phase diagram
Figure 1 shows the phase diagram on the plane of $\phi=\arctan(J_{2}/J_{1})$ and $\theta=\arctan(\Gamma_{\eta}/J_{\eta})$. Among four different regions separated by the sign of exchange interactions, we have previously studied the AF-AF region ($\phi\leq 0.5$) in Ref.[Saito and Hotta, 2024a] with a particular focus on the multicritical point observed at $\Gamma/J=\sqrt{3}$ $(\phi\approx 0.148\pi,\;\theta=\pi/3)$. This point is rigid as we find an exact solution that has degeneracy of order-$N^{2}$aito and Hotta (2024a). There are three other exact solutions(see Appendix A) , the nematic product state at $(\phi/\pi,\theta/\pi)=(0.852,0.25)$, the RVB solutionHamada et al. (1988), and decoupled ferromagnetic chain, which are linked by bold lines also partially hosting exact solutions (for nonbroken lines). These solutions are obtained systematically using the clutter-overlapped method we developed for frustration-free modelsSaito and Hotta (2024b). The point $\Gamma/J=\sqrt{3}$ is not only a multicritical but also a Lifshitz point. because the uniform ferromagnetic (F1) and the spatially modulated phase (AF-UUDD) phases meet the quantum disordered two phases, i.e. the TLL and nematic-singlet dimer phases. The low energy effective theory at around this multicritical point is shown in Appendix B, which nicely explains the Ising types of competitions among AF-UD2, UUDD, and F1 phases.
The phase diagram was determined using the DMRG methodWhite (1992, 1993) and the exact diagonalization (ED) method. For DMRG we calculate the system typically of size $N=100$ and keep up to $\chi=200$ states with up to 160 sweeps. Because the model with quantum anisotropy often exhibits orders or correlations with incommensurate or unexpected periods, we adopt the sine-square deformation (SSD)Gendiar et al. (2011) that basically suppresses the numerical biases often induced in finite-size clusters; the method that deforms the energy scale (coupling constants) of the Hamiltonian by the envelope function as ${\cal H}_{\rm ssd}=\sum_{j}h(r_{j})f(r_{j})$, using the $f(r_{j})=\sin^{2}(\pi r_{j}/(N+1))$ with $r_{j}=1,\cdots,N$. This sine-square function $f(r_{j})$ takes a maximum at the center of the system and goes to zero at both edges. It is proved both numerically and analytically that the SSD Hamiltonian offers a quantum ground state equivalent to that of the periodic boundary condition (PBC), and additionally, it has two advantages: the damping of the finite size effect and the ability to capture incommensurate orders very accurately by avoiding the bias to wavevectors commensurate with the system sizeHotta and Shibata (2012); Hotta et al. (2013). The DMRG using SSD also can evaluate the continuous magnetization curve very accurately, which is particularly useful here because the model does not conserve the total magnetization and the standard evaluation of the spin gap as the difference in the energy of total $S=0$ and $S=1$ states is not available.
II.2.1 Phase boundaries
The phase boundaries are determined numerically. The first-order transition takes place between magnetic orders of different periods and the second-order transitions are mostly the nonmagnetic-magnetic ones. We show in Fig. 2(a) the energy of the DMRG calculation for $N=100$ as a function of $\phi$. The observed kinks or the changes in the functional form provide the magnetic phase boundaries. The boundary between the nematic-singlet and the TLL phases is difficult to detect in standard methods. In Fig. 2(b) we show the $\langle S_{i}^{z}S_{i+1}^{z}-S_{i}^{x}S_{i+1}^{x}\rangle$ obtained by the Lanczos exact diagonalization method, which measures the anisotropy of nearest neighbor spin coupling. We find that the results for different $N$’s cross at the single point, which offers accurate scale-free boundary points separating TLL-ND phases.
The magnetic phase boundaries are clearly detected by the magnetization of several lattice periods and magnetization axes. Because we use the SSD, the values of magnetization are measured at the center of the system is not influenced by the boundary effect and is free of finite-size effect. Here, we measure the magnetization along the $x^{\prime}$ and $y^{\prime}$ axes which are obtained by rotating the $x$ abd $y$ axes by $\pi/4$ about the $z$-axis. Indeed, our Hamiltonian Eq.(1) remains unchanged under the $\pi$-rotation about both the $\bm{x}^{\prime}=(-1,1,0)$ and $\bm{y}^{\prime}=(1,1,0)$ axes, where the spins are transformed as $(S^{x},S^{y},S^{z})\rightarrow(-S^{y},-S^{x},-S^{z})$ and $(S^{y},S^{x},-S^{z})$, respectively. It is thus natural to consider the two as magnetic easy axes. We can thus introduce the order parameters of the F state $S^{\alpha}_{i}+S^{\alpha}_{i+1}+S^{\alpha}_{i+2}+S^{\alpha}_{i+3}$, AF-UD state $S^{\alpha}_{i}-S^{\alpha}_{i+1}+S^{\alpha}_{i+2}-S^{\alpha}_{i+3}$, and AF-UUDD state, $S^{\alpha}_{i}\pm S^{\alpha}_{i+1}-S^{\alpha}_{i+2}\mp S^{\alpha}_{i+3}$. In Fig. 3, we show these uniform and staggered magnetization along $\alpha=x^{\prime},y^{\prime}$ for two parameters across the phase diagram. In the case of $\theta=0.1\pi$, $x^{\prime}$-component of these magnetizations are exactly zero throughout $\phi$ and are not shown. Otherwise, the magnetic phases have one of the magnetizations being finite, and are exclusive to each other, capturing the phase boundaries very well. The types of magnetizations are summarized in Fig. 1(b).
In Fig. 3(c) we show the peak position of the structure factor of the two-point spin-spin correlation functions ${\cal S}^{\alpha\alpha^{\prime}}(q)=\frac{1}{N-1}\sum_{|j-j^{\prime}|=1}^{N-1}e^{iq(j-j^{\prime})}\langle S_{j}^{\alpha}S_{j^{\prime}}^{\alpha^{\prime}}\rangle$, where in the calculation we performed the SSD Fourier transformation using the envelope function Kawano and Hotta (2022). Previous studies for the $\Gamma=0$ zigzag model reported the transition from a $q=\pi$ to $q<\pi$ state of the diagonal $\alpha=\alpha^{\prime}$ structure factor inside the singlet dimer phase Hikihara et al. (2001); Bursill et al. (1995). We indeed find such a transition for $\theta>0$ ($\Gamma\neq 0$) in the present model. In particular, in the ND phase at $0.15\lesssim\phi/\pi\lesssim 0.4$, the incommensurate $q\gtrsim\pi/2$ is observed which transforms to the UUDD phase of period $q=\pi/2$ at larger $\phi$. The difference induced by the $\Gamma\neq 0$ is that the off-diagonal $xy$ component appears comparable to the diagonal ones.
Let us briefly discuss the types of symmetry breakings. As mentioned above, the spins have $x^{\prime}$ and $y^{\prime}$ as magnetic easy axis, namely the ${\mathbb{Z}}_{2}\otimes{\mathbb{Z}}_{2}$ symmetry is preseent in the Hamiltonian. In the ND and TLL phases, the spins are not ordered, namely this ${\mathbb{Z}}_{2}\otimes{\mathbb{Z}}_{2}$ is not broken. The ND phase breaks the translational symmetry but TLL does not. When the magnetic orderings take place, they break one of the ${\mathbb{Z}}_{2}\otimes{\mathbb{Z}}_{2}$ (while part of ${\mathbb{Z}}_{2}$ combined with lattice translation remains), and the spins are polarized in one of the two easy axes. The translational symmetry is broken for the AF phases, while kept for F phase. The competition among different ways of symmetry breakings generates a highly competing multicritical point.
Going back to Fig. 1(a), we see that the phase diagram has approximate reflection symmetry about the $\phi=0.5\pi$ line. By this reflection, the F1 and AF-UD1 phases are related and so as AF-UD2 and F2. The $\Gamma/J=\sqrt{3}$ tricritical point is related to another cusp point, and the TLL-singlet dimer transition point $J_{2}/J_{1}=0.2411,\Gamma_{\eta}=0$ has as the contourpart the F-Haldane dimer transition point $J_{2}/J_{1}=-1/4$, at $\Gamma_{\eta}=0$. The nearly-reflection symmetry of the phase diagram is described in the Hamiltonian as the convertion of all the spins on one of the two legs upside down. Since the spin inversion is a non-unitary transformation for $S=1/2$ it is not rigorous. Indeed, the reflection symmetry is not perfect, particularly when $\theta\lesssim 0.25\pi$. However, considering the Ising character of the magnetic phases at large $\theta$ or $\Gamma/J$, that join the low energy effective Hamiltonian near the multicritical point (see Appendix B, it is natural to find that the phase boundaries of large $\theta$ is explained very well in this context.
II.2.2 Correlation functions

### v2

@@ II.1 Model Hamiltonian
II.2 Ground-state phase diagram
II.2.1 Overview of the phase diagram
Figure 1 shows the phase diagram on the plane of $\phi=\arctan(J_{2}/J_{1})$ and $\theta=\arctan(\Gamma_{\eta}/J_{\eta})$. Among four different regions separated by the sign of exchange interactions (see the sign of $J_{1}$ and $J_{2}$ in the top panel), we have previously studied the AFM-AFM region ($\phi\leq 0.5$) in Ref.[Saito and Hotta, 2024a] with a particular focus on the multicritical point observed at $J_{2}/J_{1}=1/2,\Gamma/J=\sqrt{3}$ $(\phi\approx 0.148\pi,\;\theta=\pi/3)$. This point is rigid as we find an exact solution that has degeneracy of order-$N^{2}$ as shown in Ref.[Saito and Hotta, 2024a]. It is not only a multicritical but also a Lifshitz point, because the uniform ferromagnetic (FM1) and the antiferromagnet with up-up-down-down period (AFM-UUDD) meet the quantum disordered two phases, i.e. the TLL and nematic-singlet dimer phases. The low energy effective theory at around this multicritical point is shown in Appendix B, which nicely explains the Ising types of competitions among AFM-UD2 (antiferromagnet with up-down period), UUDD, and FM1 phases, which are summarized in Fig. 1(b). In the following subsections, we study the details of these phases.
There are three other exact solutions (see Appendix A), the nematic product state at $(\phi/\pi,\theta/\pi)=(0.852,0.25)$ ($J_{2}/J_{1}=-1/2,\Gamma/J=1$), the RVB solutionHamada et al. (1988), and decoupled ferromagnetic chain, which are linked by bold lines and also partially host exact solutions (for nonbroken lines). These states are obtained using the method we developed to have the exact MPS-based solutions for frustration-free modelsSaito and Hotta (2024b).
Here, we briefly explain how we identify the nematic product exact ground state. At $\;\Gamma_{\eta}=J_{\eta}$ and $J_{2}/J_{1}=-1/2$, the Hamiltonian has a typical frustration-free form, ${\cal H}=\sum_{l}\hat{h}_{l}$, given as the sum of operator $\hat{h}_{l}$ acting on the $l$-th triangle,
$\hat{h}_{l}=\sum_{i,j\in l}(-)^{|i-j|}J(\bm{S}_{i}\cdot\bm{S}_{j}+S^{x}_{i}S^{y}_{j}+S^{y}_{i}S^{x}_{j}),$ (3)
where we set one AFM and two FM bonds with $J\equiv J_{2}$. In a triangular unit, $\hat{h}_{l}$ has four-fold degenerate ground states of energy $-3J/4$ and the four excited states with energies ±$\pm 2(\sqrt{3}+3)J/4$. Using 0/1 representing up/down spins on $[l+1,l,l-1]$ sites on the triangle in the descending order, they are given as
$|000\rangle+i|011\rangle=|0\rangle\otimes(|00\rangle+i|11\rangle),$
$|000\rangle+i|110\rangle=(|00\rangle+i|11\rangle)\otimes|0\rangle,$ (4)
and their time reversal states, $|111\rangle-i|100\rangle$, and $|111\rangle-i|001\rangle$. The energy of the Hamiltonian consisting of $N$ triangles is $-3JN/4$ at the lowest, and the product state $|\Psi\rangle=\prod_{i=1}^{N/2}|p_{1}\rangle_{2i-1,2i}$, with $|p_{1}\rangle=(|00\rangle+i|11\rangle)/\sqrt{2}$, satisfies such energy condition, because for all choices of triangles, we find either of Eq.(4) to be the constituent. Its time reversal or translational counterparts are the other degenerate ground state. The translational symmetry is broken similarly to the Majumdar-Ghosh (MG) singlet product state at $J_{2}/J_{1}=0.5$. The other exact solutions in the phase diagram can be obtained in the same frustration-free form, although the other non-product state solutions are not explicitly written in the analytical form but rely on the MPS languageSaito and Hotta (2024b).
The phase diagram was determined using the DMRG methodWhite (1992, 1993) and the exact diagonalization (ED) method. For DMRG we calculate the system typically of size $N=100$ and keep up to $\chi=200$ states with up to 160 sweeps. Because the model with quantum anisotropy often exhibits orders or correlations with incommensurate or unexpected periods, we adopt the sine-square deformation (SSD)Gendiar et al. (2011) that suppresses the numerical biases often induced in finite-size clusters; for any given Hamiltonian ${\cal H}=\sum_{j}\hat{h}(r_{j})$ based on the local operator $\hat{h}(r_{j})$ at spatial coordinate $r_{j}=1,\cdots,N$, the method deforms its local energy scale by the envelope function as ${\cal H}_{\rm ssd}=\sum_{j}\hat{h}(r_{j})f(r_{j})$, using the $f(r_{j})=\sin^{2}(\pi r_{j}/(N+1))$. This sine-square function $f(r_{j})$ takes a maximum at the center of the system and goes to zero at both edges. It is proved both numerically and analytically that the SSD Hamiltonian offers a quantum ground state equivalent to that of the periodic boundary condition (PBC) Hikihara and Nishino (2011); Maruyama et al. (2011); Katsura (2011). Additionally, it has two advantages: the damping of the finite size effect and the ability to capture incommensurate orders very accurately by avoiding the bias to wavevectors commensurate with the system sizeHotta and Shibata (2012); Hotta et al. (2013). The boundary effects are safely excluded and the correlation functions as well as local quantities are safely evaluated, much more reliably than the open boundary onesShibata and Hotta (2011). The DMRG using SSD also can evaluate the continuous magnetization curve very accurately, which is particularly useful here because the model does not conserve the total magnetization and the standard evaluation of the spin gap as the difference in the energy of total $S=0$ and $S=1$ states is not available.
II.2.2 Phase boundaries
We first introduce the order parameters of the three magnetic phases and the two nonmagnetic phases in the diagram:
FM  $:S^{\alpha}_{i}+S^{\alpha}_{i+1}+S^{\alpha}_{i+2}+S^{\alpha}_{i+3},$
AFM-UD  $:S^{\alpha}_{i}-S^{\alpha}_{i+1}+S^{\alpha}_{i+2}-S^{\alpha}_{i+3},$
AFM-UUDD $:S^{\alpha}_{i}\pm S^{\alpha}_{i+1}-S^{\alpha}_{i+2}\mp S^{\alpha}_{i+3},$
VC  $:\kappa_{i}^{(n)}=S_{i}^{x}S_{i+n}^{y}-S_{i}^{y}S_{i+n}^{x},$
N  $:Q^{xy}_{i}=S_{i}^{x}S_{i+1}^{y}+S_{i}^{y}S_{i+1}^{x}.$ (5)
For the VC phase, we choose $\kappa_{i}^{(1)}$ which is the most susceptible to the present Hamiltonian. Regarding the nematic (N) phase, There are five independent nematic (quadrupolar) operators defined on a spin-1/2 pairs forming spin-1 given as Andreev and Grishchuk (1984)
In Eq.(1), the parameter $\Gamma$ is coupled to the $xy$ component $Q_{i}^{xy}=S_{i}^{x}S_{i+1}^{y}+S_{i}^{y}S_{i+1}^{x}$, and it works as a “field” (or a chemical potential) to condense the nematic particle represented by $Q_{i}^{xy}$, and accordingly, the other four parameters are irrelevant.
Figure 2: (a) Energy obtained by DMRG calculation and (b) $\langle S^{z}_{i}S^{z}_{i+1}-S^{x}_{i}S^{x}_{i+1}\rangle$ obtained by the exact diagonalization with PBC as functions of $\phi$. The kinks in (a) and crossings in (b) are used to determine the magnetic and nonmagnetic phase boundaries, respectively. Broken lines in panel (a) are shown as a guide to determine one of the phase boundaries.
Figure 3: Uniform and staggered magnetizations in a period of two or four sites, $S^{\alpha}_{i}+S^{\alpha}_{i+1}+S^{\alpha}_{i+2}+S^{\alpha}_{i+3}$, $S^{\alpha}_{i}-S^{\alpha}_{i+1}+S^{\alpha}_{i+2}-S^{\alpha}_{i+3}$, and $S^{\alpha}_{i}\pm S^{\alpha}_{i+1}-S^{\alpha}_{i+2}\mp S^{\alpha}_{i+3}$, for (a) $\theta/\pi=0.1$ ($\Gamma/J\approx 0.32$) in $\alpha=y^{\prime}$-direction and (b) $\theta/\pi=0.4$ ($\Gamma/J\approx 3.1$) in $x^{\prime}$ and $y$’-directions as functions of $\phi$. (c) Vector chiral correlation function $\overline{|\langle\kappa^{(n)}_{i}\kappa^{(n^{\prime})}_{j}\rangle|}$ averaged over $|i-j|=58-61$ as function $\phi$. The gray region has the relatively weak signal of chiral correlation. (d) Peak position $q$ of the structure factor ${\cal S}^{\alpha\alpha^{\prime}}(q)$ for $\alpha\alpha^{\prime}=xx,zz,xy$ as a function of $\phi$. We set $\theta/\pi=0.125$ ($\Gamma/J\approx 0.41$) for (c,d).
The phase boundaries are determined numerically. The first-order transition takes place between magnetic orders of different periods and the second-order transitions are mostly the nonmagnetic-magnetic ones. We show in Fig. 2(a) the energy of the SSD Hamiltonian of the DMRG calculation for $N=100$ as a function of $\phi$. The observed kinks or the changes in the functional form provide the magnetic phase boundaries very accurately, as has been demonsrated in previous literaturesKawano and Hotta (2023); Makuta and Hotta (2024). Namely, even if the envelope function $f(r_{j})$ is placed, which will alter the value of the total energy of the system itself, because the energy is an extensive quantity, the comparison of energies between different phases works out quantitatively as accurate as we do in PBC. We can see shortly that they coincide with the ones we derive using the magnetic order parameters. The boundary between the nematic-singlet and the TLL phases is difficult to detect in standard methods. In Fig. 2(b) we show the $\langle S_{i}^{z}S_{i+1}^{z}-S_{i}^{x}S_{i+1}^{x}\rangle$ obtained by the Lanczos exact diagonalization method, which measures the anisotropy of nearest neighbor spin coupling. We find that the results for different $N$’s cross at the single point, which offers accurate scale-free boundary points separating TLL-ND phases.
The magnetic phase boundaries are clearly detected by the magnetization of several lattice periods and magnetization axes. Because we use the SSD, the values of magnetization measured at the center of the system is not influenced by the boundary effect and is free of finite-size effects. Here, we measure the magnetization along the $x^{\prime}$ and $y^{\prime}$ axes which are obtained by rotating the $x$ and $y$ axes by $\pi/4$ about the $z$-axis. Indeed, our Hamiltonian Eq.(1) remains unchanged under the $\pi$-rotation about both the $\bm{x}^{\prime}=(-1,1,0)$ and $\bm{y}^{\prime}=(1,1,0)$ axes, where the spins are transformed as $(S^{x},S^{y},S^{z})\rightarrow(-S^{y},-S^{x},-S^{z})$ and $(S^{y},S^{x},-S^{z})$, respectively. It is thus natural to consider the two as magnetic easy axes. We can thus introduce the order parameters of the magnetic phases as in Eq.(5). In Fig. 3, we show these uniform and staggered magnetization along $\alpha=x^{\prime},y^{\prime}$ for two parameters across the phase diagram. In the case of $\theta=0.1\pi$ ($\Gamma/J\approx 0.32$), $x^{\prime}$-component of these magnetizations are exactly zero throughout $\phi$ and are not shown. Otherwise, the magnetic phases have one of the magnetizations being finite, and are exclusive to each other, capturing the phase boundaries very well. The types of magnetizations are summarized in Fig. 1(b).
In Fig. 3(d) we show the peak position of the structure factor of the two-point spin-spin correlation functions
${\cal S}^{\alpha\alpha^{\prime}}(q)=\frac{1}{N-1}\sum_{|j-j^{\prime}|=1}^{N-1}e^{iq(j-j^{\prime})}\langle S_{j}^{\alpha}S_{j^{\prime}}^{\alpha^{\prime}}\rangle,$ (7)
where in the calculation we performed the SSD Fourier transformation using the envelope function Kawano and Hotta (2022). Previous studies for the $\Gamma=0$ zigzag model reported the transition from a $q=\pi$ to $q<\pi$ state of the diagonal $\alpha=\alpha^{\prime}$ structure factor inside the singlet dimer phase Hikihara et al. (2001); Bursill et al. (1995). We indeed find such a transition for $\theta>0$ ($\Gamma\neq 0$) in the present model. In particular, in the ND phase at $0.15\lesssim\phi/\pi\lesssim 0.4$ ($0.5\lesssim J_{2}/J_{1}\lesssim 3$), the incommensurate $q\gtrsim\pi/2$ is observed which transforms to the UUDD phase of period $q=\pi/2$ at larger $\phi$. The difference induced by the $\Gamma\neq 0$ is that the off-diagonal $xy$ component appears comparable to the diagonal ones.
Let us briefly discuss the types of symmetry breakings. As mentioned above, the spins have $x^{\prime}$ and $y^{\prime}$ as magnetic easy axis, namely the ${\mathbb{Z}}_{2}\otimes{\mathbb{Z}}_{2}$ symmetry is preseent in the Hamiltonian. In the ND and TLL phases, the spins are not ordered, namely this ${\mathbb{Z}}_{2}\otimes{\mathbb{Z}}_{2}$ is not broken. The ND phase breaks the translational symmetry but TLL does not. When the magnetic orderings take place, they break one of the ${\mathbb{Z}}_{2}\otimes{\mathbb{Z}}_{2}$ (while part of ${\mathbb{Z}}_{2}$ combined with lattice translation remains), and the spins are polarized in one of the two easy axes. The translational symmetry is broken for the AFM phases, while kept for FM phase. The competition among different ways of symmetry breaking generates a highly competing multicritical point.
Going back to Fig. 1(a), we see that the phase diagram has approximate reflection symmetry about the $\phi=0.5\pi$ line. By this reflection, the FM1 and AFM-UD1 phases are related and so as AFM-UD2 and FM2. The $\Gamma/J=\sqrt{3}$ tricritical point is related to another cusp point, and the TLL-singlet dimer transition point $J_{2}/J_{1}=0.2411,\Gamma_{\eta}=0$ ($\phi=0.075\pi,\theta=0$) has as the contourpart the FM-Haldane dimer transition point $J_{2}/J_{1}=-1/4$, at $\Gamma_{\eta}=0$ ($\phi=0.922\pi,\theta=0$). The nearly-reflection symmetry of the phase diagram is described in the Hamiltonian as the conversion of all the spins on one of the two legs upside down. Since the spin inversion is a non-unitary transformation for $S=1/2$ it is not rigorous. Indeed, the reflection symmetry is not perfect, particularly when $\theta\lesssim 0.25\pi$ ($\Gamma/J\lesssim 1$). However, considering the Ising character of the magnetic phases at large $\theta$ or $\Gamma/J$, that join the low energy effective Hamiltonian near the multicritical point (see Appendix B, it is natural to find that the phase boundaries of large $\theta$ are explained very well in this context.)
@@ II.2.2 Phase boundaries
Figure 5: (a) Magnetization curves for field $h$ applied along the $z$-direction. We choose $(\phi,\theta)/\pi=(0.05,0.05)$ (($J_{2}/J_{1},\Gamma/J)\approx(0.16,0.16)$) for TLL, $(0.2,0.05)$ ((0.72,0.16)) for ND, $(0.75,0.1)$ ((-1.0,0.32)) for VC, and $(0.85,0.3)$ ((-0.51,1.4)) for N phases. Among them, only the ND phase has a spin gap $\Delta\approx 0.3$. (b) Nematic parameter $\overline{\langle Q^{xy}_{i}Q^{xy}_{j}\rangle}$ averaged over $|i-j|=58-61$ as a function of $\Gamma/J$ for the choices of $\phi/\pi$ corresponding to those of panel (a). All of them behave in power of $\Gamma/J$.
II.2.3 Correlation functions